OpenStudy (anonymous):

I need a quick answer to this please! When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 16 and 64. What is the length of the altitude? 4 8 16 32

OpenStudy (mertsj):

\[\frac{16}{a}=\frac{a}{64}\]

OpenStudy (mertsj):

\[a^2=16(64)\]

Directrix (directrix):

32

OpenStudy (mertsj):

a=32

OpenStudy (anonymous):

I got one more guys, and thanks!